OpenStudy (rjtrain1):

April wants to multiply 6 and 42 using the distributive property. Which of the following number sentences shows how she could do it? will medal

OpenStudy (rjtrain1):

6 × 42 = 42 × 6 6 × (40 + 2) = (6 × 40) + 2 (6 × 42) × 1 = 6 × (42 × 1) 6 × (40 + 2) = (6 × 40) + (6 × 2) are the options

OpenStudy (francesca_marie):

Hold on

OpenStudy (francesca_marie):

I'd go with A

OpenStudy (rjtrain1):

thank you